# SAFES: Sequential Privacy and Fairness Enhancing Data Synthesis for Responsible AI

### v1

@@ 3.2 Fairness
We separate a dataset $D$ into three disjoint sets of variables, $D=(X,G,Y)$. $Y$ is a binary response variable consisting of a “favorable” outcome (e.g., approved for a loan) and an “unfavorable” outcome. The protected attributes $G$ distinguish different groups ($g_{1},g_{2},\ldots$) (e.g., race, gender). We assume each group can be collapsed to binary attributes representing a “privileged” and an “unprivileged” group. WLOG, we assume that the favorable outcome and the privileged group are each encoded as $1$. $X$ contains the non-protected predictors. Our ultimate fairness goal is to ensure that downstream classifiers trained on a dataset are fair with respect to the protected attributes. That said, it is also useful to measure the real-world bias present in the dataset itself. We introduce several metrics for measuring the dataset bias and downstream classifier fairness.
We consider a dataset to have structural bias if the probability of an observation receiving the favorable outcome is different for observations in the privileged and unprivileged groups. This can be measured by the conditional outcome difference (COD). By definition, COD is a property of the dataset rather than of an AI or ML algorithm. When $Y$ is independent of $G$, COD = 0, which represents the least biased dataset possible by this definition.
@@
Given a dataset $D\!=\!(X,G,Y)$, let $g$ be a protected attribute in $G$. Then
@@
In statistical terms, structural bias is the same as the existence of dependency or correlation between group (e.g., privileged vs. unprivileged) and the outcome variable. Of course, not all such correlations indicate bias; some reflect genuine relationships. Here, we focus on cases where this correlation, if it exists, is not causal and would disappear if we conditioned on other relevant covariates that are not related to privilege (e.g. years of schooling). Additionally, for datasets without structural bias, imbalances in data representation between privileged and unprivileged groups can increase the likelihood of prediction bias in downstream ML algorithms. This prediction bias can be assessed using algorithmic fairness metrics, presented below.
Two commonly used fairness metrics for classifiers are statistical parity (Dwork et al.,, 2012), which ensures privileged and unprivileged groups are equally likely to get the favorable decision, and equalized odds (Hardt et al.,, 2016), which ensures privileged and unprivileged groups have identical true and false positive rates.
@@
We define the conditional utility difference (CUD) to measure the balance in the utility of a classifier between the unprivileged and privileged groups. The utility function $u$ in CUD can be defined in various ways and is context-based. For example, $u(Y,\hat{Y}|g)$ can be the classification accuracy in group $g$.
@@
CUD encompasses many fairness metrics in the literature like overall accuracy equality (Berk et al.,, 2021), predictive parity (Chouldechova,, 2017), and false positive/negative rate balance (Chouldechova,, 2017). For example, the CUD for false negative rate would use $u(Y,\hat{Y}|g)=P(\hat{Y}=0|Y=1,g)$.
For all fairness metrics presented in this section, a value closer to 00 is more fair. Negative values typically indicate unfairness in favor of the privileged group (e.g., a greater proportion of the privileged group is approved for a loan), while positive values favor the unprivileged group. We estimate each of these metrics empirically in our experiments.
4 The SAFES procedure
In this section, we first present SAFES – Sequential PrivAcy and Fairness Enhancing data Synthesis, and then briefly summarize the DP data synthesizer and the fairness-aware data transformation procedure used in our experiments, along with the reasons for selecting these methods for the experiments.
4.1 The algorithm
The SAFES algorithm consists of two sequential steps: DP data synthesis followed by fairness-aware data transformation. We provide the SAFES procedure in Algorithm 1, along with a formal claim for its DP guarantees.
Proposition 1.
SAFES satisfies DP avspacet privacy loss $\Theta_{1}$.
The proof is straightforward. Since $S$ satisfies DP with parameters $\Theta_{1}$, $D^{*}$ achieves $\Theta_{1}$-DP guarantees. Since the fairness data transformation $T$ only operates on $D^{*}$ and never accesses the original data $D$, the post-processing theorem for DP ensures the same DP guarantees on $\tilde{D}^{*}$.
Input : Dataset $D$, privacy loss parameters $\Theta_{1}$, fairness parameters $\Theta_{2}$
Output : Privacy-preserving and fairness-aware synthetic dataset $\tilde{D}^{*}$
1 Generate a privacy-preserving synthetic dataset $D^{*}$ at privacy loss $\Theta_{1}$ via a DP synthesizer $S$ (AIM in our experiments).
2 Transform $D^{*}$ with fairness parameters $\Theta_{2}$ to obtain $\tilde{D}^{*}$ via a fairness pre-processor $T$ (TOT in our experiments).
@@ 4.1 The algorithm
Algorithm 1 accommodates any DP types. For example, if $\rho$-zCDP is used, then $\Theta_{1}=\{\rho\}$; if $(\varepsilon,\delta)$-DP or $\varepsilon$-DP is used, then $\Theta_{1}=\{\varepsilon,\delta\}$ or $\Theta_{1}=\{\varepsilon\}$. Algorithm 1 also permits any legitimate DP synthesizers described in the Introduction section. Similarly, any fairness-aware data pre-processor $T$ can be applied. The immunity to post-processing property of DP permits tuning $\eta$ to achieve the desired fairness-utility trade-off without additional privacy cost, since SAFES applies privacy and fairness sequentially. Figure 1 suggests the SAFES procedure is flexible and modular, allowing users to selectively apply its steps without requiring a strict start-to-finish approach. Users can directly release the DP synthetic data if fairness is not a concern or skip the DP synthesis step to work with previously released DP synthetic data.
Figure 1: The SAFES procedure and its applications
Algorithm 1 is general to permit any DP data synthesizer and fairness data pre-processor. Our experiments use AIM (McKenna et al.,, 2022) as the DP data synthesizer and a data transformation procedure that transforms $(X,Y)\subset D$ but leaves $G$ unchanged (Calmon et al.,, 2017) as the fairness-aware transformation step, respectively. For brevity, we refer to the latter simply as Triple-cOnstrained Transformation (TOT).
We chose AIM as the DP data synthesizer in the experiments because it has been shown to consistently outperform other synthesis methods in various utility metrics in previous works on DP data synthesis (McKenna et al.,, 2022). In addition, previous studies suggest that marginal-based methods, to which AIM belong, at large privacy budgets do not exacerbate group unfairness as much as deep generative model-based DP synthesis methods (Pereira et al.,, 2024). We chose TOT over other (limited) fairness-aware pre-processing methods (Hajian and Domingo-Ferrer,, 2013) because, as pointed out in (Calmon et al.,, 2017), TOT permits explicit control over the balance between group and individual fairness and also connects well with the broader statistical learning framework due to its probabilistic transformation.
4.2 Marginal-based DP data synthesis procedure
In general, for marginal-based synthesizers, a “workload” of $k$ marginals $W=\{\boldsymbol{\mu}^{\prime}_{1},\boldsymbol{\mu}^{\prime}_{2},\ldots,\boldsymbol{\mu}^{\prime}_{k}\}$ is first specified, with each $\boldsymbol{\mu}^{\prime}_{i}$ representing a marginal on data $D$ (e.g., a 2-way marginal on race and sex). A set of $m$ marginals from $W$ are then selected (possibly with replacement) and sanitized via DP mechanisms to compute the output marginals $\hat{\boldsymbol{\mu}}=\{\hat{\boldsymbol{\mu}}_{1},\ldots,\hat{\boldsymbol{\mu}}_{m}\}$. From the estimated DP marginal counts, a distribution $\hat{p}$ is learned that minimizes the difference between its marginals and $\hat{\boldsymbol{\mu}}$. $\hat{p}$ can then be sampled to generate synthetic data $D^{*}$ with DP guarantees.
AIM, the method used for the DP data synthesis subroutine of the SAFES procedure in our experiments, makes several improvements to this standard procedure, as listed in Algorithm 2.
@@ 4.2 Marginal-based DP data synthesis procedure
Input : Dataset $D=(X,G,Y)$ with $d$ variables, marginal workload $W=\{\boldsymbol{\mu}^{\prime}_{1},\ldots,\boldsymbol{\mu}^{\prime}_{k}\}$, privacy loss $\rho$
@@
Initialize $\sigma_{0}$, $\xi_{0}$ // conservative initialization is recommended
@@
2 $\hat{p}_{0}\leftarrow\arg\min_{p}\sum_{i=1}^{d}\left\lVert\boldsymbol{\mu}_{0,p}[i]-\hat{\boldsymbol{\mu}}_{0}[i]\right\rVert_{2}^{2}/\sigma_{0}$, where $\boldsymbol{\mu}_{0,p}$ contains all 1-way marginals based on distribution $p$ for data $D$
@@ 3 $\rho_{\textnormal{used}}\leftarrow d/(2\sigma^{2}_{0})$
for $j=1$ to $k$ do $w_{i}\leftarrow\sum_{i=1}^{k}|\boldsymbol{\mu}^{\prime}_{j}\cap\boldsymbol{\mu}^{\prime}_{i}|$ end // ||$\cdot$|| ⋅ | is the cardinality
@@ 4 $t\leftarrow 0$; $\sigma_{t+1}\leftarrow\sigma_{0}$; $\xi_{t+1}\leftarrow\xi_{0}$
5 while $\rho_{\textnormal{used}}<\rho$ do
@@ 6       $t\leftarrow t+1$
      Update $\sigma_{t+1}$ and $\xi_{t+1}$ // according to Algorithm 3 of AIM arXiv paper (McKenna et al.,, 2022)
8
@@
Algorithm 2 the AIM procedure for $\rho$-zCDP guarantees(McKenna et al.,, 2022)
For illustration, the steps in Algorithm 2 are listed with $\rho$-zCDP guarantees, but other DP types with the associated mechanisms can be used. The user first provides a workload $W$ to be considered for learning/updating a DP graphical model. The initial graphical model $\hat{p}_{0}$ is formed from all one-way marginals, then adaptively updated with higher-order marginals selected with replacement from $W$ in order to optimize the utility of the graphical model. The privacy budget spending scheme and update rule ($\sigma_{t+1}$ and $\xi_{t+1}$ on line 13 of Algorithm 2) ensure that marginals measured in later iterations receive a greater portion of the privacy budget when necessary to ensure useful marginal measurements. In particular, if the difference between the marginals produced by $\hat{p}_{i}$ and $\hat{p}_{i-1}$ is small, meaning little information is gained this round, $\xi_{i+1}=2\xi_{i}$ and $\sigma_{i+1}=\sigma_{i}/2$. See Algorithm 3 in the arXiv version of the AIM paper (McKenna et al.,, 2022) for more details. Considerations are also taken to ensure computational tractability for the graphical representation. These iterations proceed until the privacy budget is exhausted, at which point the current graphical model is sampled to obtain the synthetic dataset.
Since AIM only works with categorical or ordinal variables, for continuous variables, some discretization will be needed before applying AIM.
4.3 Fairness-aware data transformation
The TOT method transforms $(X,Y)$ to $(\tilde{X},\tilde{Y})$, but leaves $G$ unchanged. We summarize the main idea of TOT below and refer readers to the original paper (Calmon et al.,, 2017) for more details and theoretical results. TOT learns a randomized mapping $T$ in the form of a conditional distribution $q_{\tilde{X},\tilde{Y}|X,G,Y}$ by solving an optimization problem that balances controlling discrimination between groups (group fairness), limiting distortion of individual observations (individual fairness), and maintaining a data distribution similar to the one before the transformation (utility preservation). For discrimination control, it requires $q_{\tilde{Y}|G}$ be similar for any two groups. Specifically,

### v3

@@ 2.2 Fairness
This work primarily focuses on binary outcome variables, as this is the primary setting for most fairness literature and metrics Mehrabi et al., (2021) and lends itself nicely to many real-world scenarios where fairness is a concern such as hiring and loan decisions. We seek to improve group fairness, meaning we desire that the privileged group does not receive undue over-representation among those with favorable outcomes/predictions based on group status.
We separate a dataset $D$ into three disjoint sets of variables, $D=(X,G,Y)$. $Y$ is a binary variable consisting of “favorable” (e.g., approved for a loan) and “unfavorable” outcomes. The protected attributes $G$ distinguish different groups ($g_{1},g_{2},\ldots$) (e.g., race, gender). We assume each group variable can be collapsed to a binary attribute representing “privileged” vs “unprivileged”. WLOG, we assume that the favorable outcome and the privileged group are each encoded as $1$. $X$ contains the non-protected predictors. To ensure that downstream classifiers trained on a dataset are fair with respect to the protected attributes, it is important to also assess the bias present in the dataset itself. We introduce several metrics for measuring the dataset bias and downstream classifier fairness.
We consider a dataset to have structural bias if the probability of an observation receiving the favorable outcome is different for observations in the privileged and unprivileged groups. In statistical terms, structural bias is equivalent to non-zero correlation between group and the outcome variable. Of course, not all such correlations indicate bias; some reflect genuine relationships. Here, we focus on cases where this correlation, if it exists, is assumed to be non-causal and would disappear if we conditioned on other relevant covariates that are not related to privilege (e.g. years of schooling). We introduce conditional outcome difference in Definition 6 to measure the structural bias.
@@
Let $g$ be a protected attribute in $G$ in dataset D=(X,G,Y)D\!=\!(X,G,Y), then
@@
By definition, COD is a property of the dataset rather than of an AI or ML algorithm. When $Y$ is independent of $G$, COD = 0, which represents the least bias possible based on this definition.
For datasets without structural bias, imbalances in data representation between privileged and unprivileged groups can increase the likelihood of prediction bias in downstream ML algorithms, which can be assessed using algorithmic fairness metrics. Two such commonly used metrics for classifiers are statistical parity (Dwork et al.,, 2012), which ensures privileged and unprivileged groups are equally likely to get the favorable decision, and equalized odds (Hardt et al.,, 2016), which ensures privileged and unprivileged groups have identical true and false positive rates.
@@
We additionally define the conditional utility difference to measure the balance in the utility of a classifier between the unprivileged and privileged groups. The utility function $u$ in CUD can be defined in various ways and is context-based. For example, $u(Y,\hat{Y}|g)$ can be the classification accuracy in group $g$ or false negative rate $P(\hat{Y}=0|Y=1,g)$.
@@
With the flexibility and generality of $u$, CUD encompasses many existing fairness metrics, such as overall accuracy equality (Berk et al.,, 2021), predictive parity (Chouldechova,, 2017), and false positive/negative rate balance (Chouldechova,, 2017).
For all fairness metrics presented in this section, a value closer to 0 is more fair. Negative values typically indicate unfairness in favor of the privileged group (e.g., a greater proportion of the privileged group is approved for a loan), while positive values favor the unprivileged group. We estimate each of these metrics empirically in our experiments.
3 Approach
3.1 The SAFES procedure
SAFES stands for Sequential PrivAcy and Fairness Enhancing data Synthesis. It is comprised of a DP data synthesizer and a fairness-aware data processor. Fig. 1 provides a visualization of SAFES. The figure suggests that SAFES is flexible and modular, allowing users to selectively apply its steps without requiring a strict start-to-finish approach. Users can also directly release the DP synthetic data if fairness is not a concern or skip the DP synthesis step to work with previously released DP synthetic data.
Figure 1: The SAFES procedure and its applications
Accordingly, the SAFES algorithm, as presented in Algorithm 1, consists of two sequential steps: DP data synthesis followed by fairness-aware data preprocessing. The formal claim for its DP guarantees is given in Proposition 9.
Input : Dataset $D$, DP privacy loss parameters $\Theta_{1}$, fairness parameters $\Theta_{2}$ if applicable.
Output : Privacy-preserving and fairness-aware synthetic dataset $\tilde{D}^{*}$.
1
21exGenerate a privacy-preserving synthetic dataset $D^{*}$ at privacy loss $\Theta_{1}$ via a DP synthesizer $S$.
3 Transform $D^{*}$ (with fairness parameters $\Theta_{2}$ if applicable) to obtain $\tilde{D}^{*}$ via a fairness data processor $T$.
@@ 3.1 The SAFES procedure
Proposition 9.
SAFES satisfies DP at privacy loss $\Theta_{1}$.
The proof is straightforward. Since $S$ satisfies DP with parameters $\Theta_{1}$, $D^{*}$ achieves $\Theta_{1}$-DP guarantees. Since the fairness data preprocessing $T$ only operates on $D^{*}$ and never accesses the original data $D$, the post-processing theorem for DP ensures the same DP guarantees on $\tilde{D}^{*}$.
Algorithm 1 accommodates any DP types adopted by a DP synthesizer. For example, if $\rho$-zCDP is used, then $\Theta_{1}=\{\rho\}$; if $(\varepsilon,\delta)$-DP or $\varepsilon$-DP is used, then $\Theta_{1}=\{\varepsilon,\delta\}$ or $\Theta_{1}=\{\varepsilon\}$. The immunity to post-processing property of DP allows examining and comparing different fairness parameters $\Theta_{2}$, if a fairness preprocessing approach involves such a parameter, to achieve a desired fairness-utility trade-off at no additional privacy cost, a a key advantage of SAFES’s sequential treatment of privacy and fairness.
3.2 DP data synthesizer
Algorithm 1 is sufficiently general to permit any DP data synthesizer. We list two such procedures that are used in our experiments below. The first is the Adaptive and Iterative Mechanism (AIM) procedure (McKenna et al., 2022a, ) and the second is Differentially Private Conditional Tabular Generative Adversarial Network (DP-CTGAN) (Fang et al.,, 2022). AIM can provide strong utility for generated DP synthetic data but it only works for data with categorical or ordinal variables; if a dataset contains continuous variables, discretization will be needed before applying AIM, resulting in information loss. DP-CTGAN does not have such restriction and can work with both continuous and discrete variables.
The AIM procedure
AIM has been shown to consistently outperform other synthesis methods in various utility metrics in previous works (McKenna et al., 2022a, ). In addition, previous studies suggest that marginal-based methods, to which AIM belong, for large privacy loss do not exacerbate group unfairness as much as deep generative model-based DP synthesis methods (Pereira et al.,, 2024). In general, for marginal-based synthesizers, a workload $W=\{\boldsymbol{\mu}^{\prime}_{1},\boldsymbol{\mu}^{\prime}_{2},\ldots,\boldsymbol{\mu}^{\prime}_{k}\}$ is first specified, where each $\boldsymbol{\mu}^{\prime}_{i}$ for $i=1,\ldots,k$ represents a marginal of soem order on data $D$. The initial data distribution estimate $\hat{p}$ (referred to as graphical model in (McKenna et al., 2022a, )) is formed that minimizes the difference between its one-way marginals and the DP one-way marginals calculated from data $D$. $\hat{p}$ is then adaptively updated with higher-order marginals given a set of marginals selected from $W$ (possibly with replacement) with DP guarantees. When all the pre-specified privacy budget is fully used, the iteration stops and the most updated $\hat{p}$ is sampled to generate synthetic data $D^{*}$ to release.
The pseudo-code for AIM is illustrated in Algorithm 2 with $(\epsilon,\delta)$-DP and $\rho$-zCDP guarantees, but other DP types with the associated mechanisms can be used. The privacy budget spending scheme and update rule ($\sigma_{t+1}$ and $\xi_{t+1}$ on line 13 of Algorithm 2) ensure that marginals measured in later iterations receive a greater portion of the privacy budget when necessary to ensure useful marginals are learned in each iteration $t$. In particular, if the difference between the marginals produced by $\hat{p}_{t}$ and $\hat{p}_{t-1}$ is small, meaning little information is gained in iteration $t$, $\xi_{t+1}=2\xi_{t}$ and $\sigma_{t+1}=\sigma_{t}/2$. See Algorithm 3 in the arXiv version of the AIM paper (McKenna et al., 2022b, ) for more details. Considerations are also taken to ensure computational tractability for the graphical representation. These iterations proceed until the privacy budget is exhausted, at which point the current graphical model is sampled to obtain the synthetic dataset.
The DP-CTGAN synthesizer
DP-CTGAN builds on CTGAN’s conditional sampling and mode-specific normalization for continuous columns with DP guarantees through DP-SGD. Algorithm 3 summarizes the training steps. There are $m$ conditional draws given a particular level of a categorical variable within each iteration. For each $j\in\{1,\dots,m\}$, a categorical mask $M_{j}$ and conditional vector $\texttt{cond}_{j}$ are sampled to balance the discrete levels; the GAN generator produces a fake sample $\hat{r}_{j}=G_{\Phi_{\mathcal{G}}}(z_{j},\texttt{cond}_{j})$ given a sample $z_{j}\!\sim\!\mathcal{N}(0,I)$, and the GAN critic is refined with $s$ condition-matched real–vs–synthetic comparisons via a conditional loss. The generator is then updated with a loss function that combines a loss term aligned with $M_{j}$ for $j=1,\dots,m$ and an adversarial term against the current critic, and an optional gradient penalty term. Iterations continue until the pre-specified privacy budget is exhausted, after which the trained generator generates the synthetic dataset $D^{*}$.
@@ 3.2 DP data synthesizer
Input : Tabular data D=(X,G,Y)D\!=\!(X,G,Y) with $d$ variables, marginal workload $W\!=\!\{\boldsymbol{\mu}^{\prime}_{1},\ldots,\boldsymbol{\mu}^{\prime}_{k}\}$, privacy budget $(\varepsilon_{0},\delta_{0})\!\!$
@@
1
1exInitialize $\sigma_{0}$, $\xi_{0}$ // conservative initialization is recommended
2 Convert $(\varepsilon_{0},\delta_{0})$ to target $\rho_{0}$ in zCDP using the conversion theorem 3
@@
4 $\hat{p}_{0}\leftarrow\arg\min_{p}\sum_{i=1}^{d}\left\lVert\boldsymbol{\mu}_{0,p}[i]-\hat{\boldsymbol{\mu}}_{0}[i]\right\rVert_{2}^{2}/\sigma_{0}$, where $p$ denotes the unknown underlying distribution for data $D$ and $\boldsymbol{\mu}_{0,p}$ contains all 1-way marginals given $p$
@@ 5 $\rho_{\textnormal{used}}\leftarrow d/(2\sigma^{2}_{0})$
$w_{j}\leftarrow\sum_{i=1}^{k}|\boldsymbol{\mu}^{\prime}_{j}\cap\boldsymbol{\mu}^{\prime}_{i}|$ for $j=1$ to $k$ // $|\cdot|$ is the cardinality
@@ 6 $t\leftarrow 0$; $\sigma_{t+1}\leftarrow\sigma_{0}$; $\xi_{t+1}\leftarrow\xi_{0}$
7 while $\rho_{\textnormal{used}}<\rho_{0}$ do
@@ 8 $t\leftarrow t+1$
  Update $\sigma_{t+1}$ and $\xi_{t+1}$ // according to Algorithm 3 of AIM arXiv paper (McKenna et al., 2022b, )
10
@@
Algorithm 2 The AIM procedure (McKenna et al., 2022a, )
Input : Tabular dataset $D$ with discrete columns $\mathbf{d}_{c}$, batch size $m$, critic steps $s$, learning rate $\alpha$, clipping bound $b$, noise scale $\sigma$, gradient-penalty weight $\lambda$, target privacy budget $(\varepsilon_{0},\delta_{0})$, privacy accountant $\mathcal{A}$
Output : Privacy-preserving synthetic dataset $D^{*}$
1
21exPreprocess $D$ (one-hot for discrete variables, mode-specific normalization for continuous variables)
3 Initialize parameters $\Phi_{\mathcal{G}}$ and $\Phi_{\mathcal{C}}$ for generator $\mathcal{G}$ and critic $\mathcal{C}$
4 $\varepsilon_{\textnormal{used}}\leftarrow 0$
5 while $\varepsilon_{\textnormal{used}}<\varepsilon_{0}$ do
6 for $j=1$ to $m$ do
7      Construct conditional categorical mask $M_{j}$ over $\mathbf{d}$ and sample conditional vector $\texttt{cond}_{j}$
8      Sample $z_{j}\sim\mathcal{N}(\mathbf{0},I)$ and generate fake row $\hat{r}_{j}\leftarrow\mathcal{G}_{\Phi_{\mathcal{G}}}(z_{j},\texttt{cond}_{j})$
9      Sample real row $r_{j}\sim D\,|\,\texttt{cond}_{j}$
10
11     Sample $\texttt{cond}_{k}^{(j)}$, $r_{k}^{(j)}$, and $\hat{r}_{k}^{(j)}$ at $k=1,\ldots,s$
12      Compute critic loss $L_{\mathcal{C}}^{(j)}\;\leftarrow\;s^{-1}\sum_{k=1}^{s}\!\Big(\mathcal{C}_{\Phi_{\mathcal{C}}}(\hat{r}_{k}^{(j)},\texttt{cond}_{k}^{(j)})-\mathcal{C}_{\Phi_{\mathcal{C}}}(r_{k}^{(j)},\texttt{cond}_{k}^{(j)})\Big)\,$ .
13      Draw DP noise $\mathbf{e}_{j}\sim\mathcal{N}(\mathbf{0},(\sigma b)^{2}I)$;
      Update $\Phi_{\mathcal{C}}\leftarrow\Phi_{\mathcal{C}}-\alpha\,\texttt{Adam}\Big(\nabla_{\Phi_{\mathcal{C}}}\big(L_{\mathcal{C}}^{(j)}\;+\;\lambda\,L_{\mathrm{GP}}\big)\;+\mathbf{e}_{j}\Big)$
    // $L_{\mathrm{GP}}$ is the optional gradient penalty term
14
15   end for
16  Compute generator loss $L_{\mathcal{G}}\;\leftarrow\;m^{-1}\sum_{j=1}^{m}\mathrm{CrossEntropy}(\hat{\mathbf{d}}_{c}[j],M_{j})\;-\;s^{-1}\sum_{k=1}^{s}C_{\Phi_{\mathcal{C}}}\!\big(\hat{r}^{(s)}_{k},\texttt{cond}^{(s)}_{k}\big)$
17   Update $\Phi_{\mathcal{G}}\leftarrow\Phi_{\mathcal{G}}-\alpha\,\texttt{Adam}(\nabla_{\Phi_{\mathcal{G}}}L_{\mathcal{G}})$
   Update $\varepsilon_{\textnormal{used}}\leftarrow\mathcal{A}\!\left(\varepsilon_{\textnormal{used}},\,\delta_{0};\,m,b,\sigma\right)$
 // cumulative $(\varepsilon,\delta)$ via the accountant
18
19 end while
20Sample $D^{*}$ from $G_{\Phi_{\mathcal{G}}}$
21 return $D^{*}$
Algorithm 3 The DP-CTGAN procedure with for $(\varepsilon,\delta)$-DP guarantees (Fang et al.,, 2022)
3.3 Fairness-aware data preprocessing
Similar to DP synthesizers, SAFES is general to permit any fairness data preprocessing procedures. There are limited fairness-aware data preprocessing methods in the literature. We use two such procedures in our experiments. The first transforms $(X,Y)\subset D$ but leaves $G$ unchanged (Calmon et al.,, 2017). The loss function used in procedure has to satisfy three constrains, as detailed later; for that reason and for convenience, we refer to the procedure as Triple-constrained Transformation (TOT). The second is a reweighting (RW) procedure (Kamiran and Calders,, 2012) that keeps $(X,Y,G)$ unchanged but assigns instance weights so the weighted empirical distribution makes $G$ and $Y$ independent for downstream training or learning with $D$.
The TOT procedure
[…]